James Company purchased a new machine that had an initial cost of $58,000. The estimated life of the machine is six years with no disposal value. If James planned on investing an additional $20,000 at the end of the sixth year, the machine would last an additional four years beyond the time expected. The effective tax rate for James Company is 40-percent and extending the life of the machine will generate a gross margin of $4,000 in each of the additional years. Assuming a discount rate of 6-percent, determine the net benefit (or loss) from investing the additional $20,000 at the end of the sixth year.
Answer #1

Net cash inflow per year = 4000- 40% = 2400

net loss = 8700 - 20000

= 11300Present hawe of future Cash Inflows. Here C = 2400 :41:4/100 2 0.04 n= u years. By solving above: 2400 - 610.044 0.04 & Quoot